OpenStudy (anonymous):

Factor: 4x2 + 12x + 9 Answer (2x + 3)2 (4x + 9)(x + 1) (2x + 1)(2x + 9) (2x + 3)(2x - 3)

Directrix (directrix):

You may recognize the given expression as a perfect square trinomial and use that pattern. if not, the factor by grouping. Look for two numbers that multiply to 36 and add to 12: 4x2 + 12x + 9 = 4x2 + 6x + 6x + 9 = 2x (2x +3) + 3(2x + 3) = (2x + 3 ) ( ? + ?) ---------------> @charlieberzak